OpenStudy (anonymous):

Match each item in Column A to an answer in Column B. 2. (1 pt) What is the distance between the two points on a number line? Column A Column B 1. X = –6, Y = 11 2. P = 8, Q = –15 3. U = –3, V = –20 4. J = 16, K = 7 A. 9 B. 17 C. 23 Question

OpenStudy (kitten_is_back):

omg K12 RILEASE SCREEN SHOT THIS sorrry caps were on

OpenStudy (anonymous):

how do you do that

OpenStudy (kitten_is_back):

do yuo have lightshot?

OpenStudy (anonymous):

no

OpenStudy (kitten_is_back):

got make a new tab and go to lightshot download it should be free and after downloading it should show a kitty

OpenStudy (anonymous):

https://app.prntscr.com/en/index.html

OpenStudy (anonymous):

@JoeDeWise

OpenStudy (anonymous):

@Sushi121212

OpenStudy (anonymous):

I do not know. :/

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

thx

OpenStudy (sushi121212):

To find the difference of a number, just subtract: 11-(-6) or 11+6 = 17 8-(-15) or 8+15 = 23 -3-(-20) or -3+20 = 17 16-7 =9

OpenStudy (sushi121212):

Did it help?

OpenStudy (anonymous):

not realy

OpenStudy (sushi121212):

The answer to #1 is 17 because in order to find the difference (or distance) of 2 numbers, just subtract, so I subtracted 11 by -6. A negative number and a subtraction sign makes it a positive so: 11-(-6) or 11+6 which equals 17. The same goes for the other questions: 8-(-15) or 8+15 = 23 -3-(-20) or -3+20 = 17 16-7 =9

OpenStudy (sushi121212):

Is this a better explanation?